OpenStudy (anonymous):

integrate (x+10)/(2x^2+5x-3)

OpenStudy (mr.math):

Do you know how to do partial fractions?

OpenStudy (mr.math):

Try to find A and B such that: \[\frac{x+10}{(x+3)(2x-1)}=\frac{A}{x+3}+\frac{B}{2x-1}.\]

OpenStudy (mr.math):

Then the integration will be trivial.